write an equation and solve the question.find the measure of an angle that is twice as large as its supplement
Answer by checkley77(12844)

2X+X=180
3X=180
X=180/3
X=60 FOR THE SMALLER ANGLE
180-60=120 FOR THE LARGER ANGLE.